. A procurement manager has been tracking delivery performance of her suppliers. Based on the following data

1. Calculate sigma performance of these suppliers using Z Score tables and rank order them

2. Comment on the performance of these suppliers

Supplier A B C
Number of deliveries 25 54 70
Number of late deliveries 3 10 12
Answer #1

(1) Answer :-

Find out the confidence levle of each supplier we can assume that the foll are the confidence level.

Number of successful deliveries for A = Number of deliveries - Number of late deliveries

Number of successful deliveries for A = 25 - 3 = 22

Number of successful deliveries for B = 54 - 0 = 44

Number of successful deliveries for C = 70 - 12 = 58.

Confidence level can be calculated using the formulae = (Number of successful delivery/total number of delivery)

For A = 22/25 = 0.88

For B = 44/54 = 0.81

For C = 58/70 = 0.82

For these we find the z values corresponding for each supplier.

oach 8 Level clciveites F& A Fa B Fo C 그2 = o.88 25 E S8 ②.samou.aepaese d. s-Btomttemaan oud.nthis S6 34 7 .3)